A carpenter used 44 ft of molding in three pieces to trim a garage door. if the long piece was 4 ft longer than three times the length of each shorter piece, then how long was each piece?
Found 2 solutions by macston, TimothyLamb:
Answer by macston(5194)

shorter pieces: length=L
longer piece: 3L+4 ft
total length=44 ft=2short pieces (2L)+ long piece(3L+4)
44 ft=2L+3L+4=5L+4 subtract 4 from each side
44-4 ft=5L+4-4
40 ft=5L divide each side by 5
40/5 ft=5L/5
8 ft=L The shorter pieces are 8 ft long
3L+4=3(8)+4=28 ft The longer piece is 28 ft
CHECK:
Short + short + long= 44 ft
8 ft + 8 ft + 28 ft=44 ft
44 ft=44 ft

Answer by TimothyLamb(4379)

x = length of the long piece (feet)
y = length of each of the two short piece (feet)
---
x + 2y = 44
x = 3y + 4
---
put the system of linear equations into standard form
---
x + 2y = 44
x - 3y = 4
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 28
y = 8
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---